Differential Equations class, please give general solution

1 2 6. x(t) = 1 1 -5 1 -1 3 X(t) -8
Answer #1

# Given x lt) - XIV) - 2 -8 -5 -3 1-1 2 I-d -5 -3-1 -) X1H) = AXTL) Let eigenvalue of A is then (A-I = 0 - 1-8 - (+1)(4+2)